# Descrip3D: Enhancing Large Language Model-based 3D Scene Understanding with Object-Level Text Descriptions

### v1

@@ Abstract
Understanding 3D scenes goes beyond simply recognizing objects; it requires reasoning about the spatial and semantic relationships between them. Current 3D scene-language models often struggle with this relational understanding, particularly when visual embeddings alone do not adequately convey the roles and interactions of objects. In this paper, we introduce Descrip3D, a novel and powerful framework that explicitly encodes the relationships between objects using natural language. Unlike previous methods that rely only on 2D and 3D embeddings, Descrip3D enhances each object with a textual description that captures both its intrinsic attributes and contextual relationships. These relational cues are incorporated into the model through a dual-level integration: embedding fusion and prompt-level injection. This allows for unified reasoning across various tasks such as grounding, captioning, and question answering, all without the need for task-specific heads or additional supervision. When evaluated on five benchmark datasets, including ScanRefer, Multi3DRefer, ScanQA, SQA3D, and Scan2Cap, Descrip3D consistently outperforms strong baseline models, demonstrating the effectiveness of language-guided relational representation for understanding complex indoor scenes.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly transformed human-computer interaction by equipping machines with unprecedented capabilities in language understanding, reasoning, and open-ended dialogue. These models have demonstrated remarkable success in a wide range of domains, including information retrieval, code generation, and multi-modal learning [17, 35, 60, 28, 49, 42]. Building on this progress, researchers have begun to extend LLMs to 3D scene understanding [33, 57, 37], aiming to empower models with the ability to interpret and reason about complex visual and spatial contexts within real-world environments.
In particular, indoor scenes present unique challenges for 3D scene-language modeling, including dense layouts, ambiguous viewpoints, and complex spatial relationships. Earlier methods [4, 11] were typically designed for specific tasks such as grounding or captioning. In contrast, recent efforts [20, 22] explore 3D multimodal large language models (MLLMs) that integrate language with point clouds and multiview images to jointly support a wide range of tasks, such as 3D visual grounding, dense captioning, and question answering, using a single trained model. Models like Chat-Scene [22] adopt object-centric pipelines by associating 2D/3D embeddings with object identifiers and feeding them into the language model. However, these approaches still show limited performance on tasks requiring spatial or semantic reasoning across multiple objects, primarily because they lack explicit modeling of inter-object relationships and underutilize the language understanding capabilities of LLMs. Although recent graph-based methods [56] attempt to address this, they introduce additional training complexity and still fall short in performance.
@@
This dual-level integration, at both the representation and the language interface levels, allows the model to leverage the prior knowledge of the LLM about the spatial and functional relations learned from the language. Crucially, this approach does not introduce architectural changes, does not introduce task-specific heads, and does not require additional human annotation. It is modular, generalizable, and easy to integrate into existing object-centric pipelines.
Empirically, our method yields consistent gains across five standard 3D scene language benchmarks. ScanRefer [4], Multi3DRefer [58], Scan2Cap [11], ScanQA [2], and SQA3D [34]. Improvements are especially pronounced in tasks involving complex grounding or multi-object reasoning, highlighting the benefit of explicit text-based relational cues in enhancing 3D scene understanding.
@@ 2 Related work
2.1 3D Vision-Language Understanding
3D vision-language understanding combines 3D spatial perception with natural language to interpret and interact with complex environments. Earlier work focused on single-task models for visual grounding [30], dense captioning [53], and question answering [29]. For example, models such as ScanRefer [4], ReferIt3D [1], and Multi3DRefer [58] focus on grounding by locating target objects based on descriptions of natural language. In terms of dense captioning, models such as the 3DVG-Transformer [59] and Vote2Cap-DETR++ [9] work to both localize and describe objects within a 3D environment. Furthermore, question-answer models like ScanQA [2] and SQA3D [34] extend the task space by incorporating a broader understanding of the scene to answer general or context-dependent questions.
Recent research is increasingly focusing on multi-task frameworks that unify these objectives to enhance representation sharing and robust performance. For example, methods such as 3DJCG [3] and D3Net [5] unify 3D visual grounding and dense captioning through shared encoders and task-specific heads, leveraging complementary supervision across tasks to improve performance. Likewise, 3D-VisTA [62] and 3D-VLP [27] employ large-scale pre-training to learn generalizable 3D vision-language representations, enhancing performance across diverse downstream tasks. However, their reliance on task-specific heads and visual embeddings treats language as auxiliary, lacking explicit modeling of semantic object relationships.
2.2 Multi-modal Large Language Models
LLMs have demonstrated impressive capabilities in reasoning, generalization, and multi-turn dialogue, sparking interest in extending their input space to include 3D data [33]. In the 3D vision language domain, models such as PointLLM [48], ImageBind-LLM [18], Point-Bind [16], Ulip [50], and CG3D [19] extend LLMs to handle 3D data by aligning the embeddings of the point cloud with the language embedding space. These methods enable object-level understanding through joint representations of geometric and linguistic information, facilitating tasks such as 3D object captioning and question answering. However, they primarily operate on isolated object input and lack mechanisms for encoding explicit relationships between objects, which are essential for capturing scene-level semantics.
Multimodal large language models (MLLMs) have recently been extended to 3D scene understanding to support open-ended tasks such as question answering. Approaches such as 3D-LLM [20] map the embeddings of point clouds into the language space using position embeddings, allowing basic multimodal reasoning but lacking fine-grained localization and interactive grounding. Chat-3D [44] enhances the 3D dialogue by using object-centric prompts and a three-stage training scheme to align the characteristics and relations of objects with language models. However, its architecture restricts the model’s attention to individually selected objects, limiting its ability to reason in broader contexts of the scene. Chat-Scene [22] further advanced this line by incorporating object identifiers and aligning them with 2D and 3D visual embeddings, enabling strong performance on scene-level tasks such as question answering and visual grounding. Yet, they still lack explicit mechanisms for modeling inter-object relationships, limiting relational and spatial reasoning.
@@ 2.3 Explicit Object Relationship Representation
Although prior models lack explicit mechanisms for encoding inter-object relationships, recent efforts have begun to address this limitation through graph-based or language-based modeling. ConceptGraph [15] constructs scene-level concept graphs by aligning 3D object embeddings with textual descriptions using vision-language models, enabling semantic relationship modeling and open-vocabulary generalization. HOV-SG [45] advances this direction by generating holistic object-scene graphs from 3D scans, capturing spatial and semantic relations across the scene. However, such graph-based representations often require long token sequences and suffer from high inference costs in large-scale scenes. To address this, 3DGraphLLM [56] incorporates knowledge graph embeddings into object representations, encoding pairwise relationships to support relational reasoning. Yet, this approach introduces significant combinatorial overhead by enumerating object-object triplets and suffers from error propagation in graph construction. Notably, the addition of explicit graph structures in 3DGraphLLM leads to a performance drop on downstream question answering tasks, suggesting that current graph-based formulations may hinder rather than help relational reasoning in practice. Scene-LLM [14] instead leverages LLMs with scene-level prompts to generate global summaries and captions, but lacks fine-grained object-level relational modeling. Thus, explicitly and efficiently representing object relationships in 3D scenes remains an open and challenging problem.
@@ 3.1 Overview
Our method empowers LLMs to perform precise and context-aware reasoning over complex 3D scenes by augmenting object representations with rich, relational textual descriptions. We introduce a powerful text-based relational modality that explicitly encodes both spatial and semantic relationships between objects. Our framework integrates pre-trained 3D and 2D visual encoders, a text encoder, and an LLM, and represents each object as a fused token embedding that combines geometric structure, visual appearance, and relational context. These relational cues are injected into both the object-level embeddings and directly into the LLM prompt, enabling fine-grained multimodal understanding and interaction. This dual-level integration significantly improves the ability of LLM to reason about object relationships, supporting downstream tasks such as grounding, captioning, and question answering with high accuracy and interpretability. An overview of our pipeline is shown in Figure 2.
@@ 3.2 Object-Centric Scene Representation
We decompose each scene into discrete object proposals using Mask3D [39] first, producing a set of segmented object point clouds $\{\mathbf{P}_{1},...,\mathbf{P}_{n}\}$. Each object $\mathbf{P}_{i}\in\mathbb{R}^{m_{i}\times 6}$ includes XYZ coordinates and RGB values. Here, $m_{i}$ refers to the number of points in the proposal of the $i$th object. Each resulting object is treated as a fundamental unit for understanding the scene. To support unambiguous reference and interaction, we adopt the design introduced in Chat-Scene [22] to associate each object with a unique learnable identifier token $<\!OBJ_{i}\!>$. These tokens are integrated into the tokenizer vocabulary and embedded alongside the embeddings of the object.
@@ 3.3 Visual Embedding Extraction
For each detected object, we extract geometric and visual embeddings from 3D point clouds and multi-view images. To capture 3D geometry, we use the pre-trained Uni3D [61] encoder, which provides high-quality spatial embeddings by processing each object’s segmented point cloud $\mathbf{P}_{i}$, producing an embedding vector $\mathbf{Z}^{v_{p}}_{i}\in\mathbb{R}^{1\times d}$. Uni3D is chosen for its strong performance in capturing fine-grained geometric structure across diverse 3D scenes. For 2D appearance, we employ the pre-trained DINOv2 [36] vision transformer to extract embeddings from multi-view images. The 3D mask of each object is projected onto these images, and the corresponding regions are cropped from the embedding maps and aggregated into a 2D visual embedding $\mathbf{Z}^{v_{f}}_{i}\in\mathbb{R}^{1\times d}$. DINOv2 is used for its ability to extract semantically rich embeddings from high-resolution visual input.
@@ 3.4 Text-Based Relational Modality
While the embeddings extracted from both 2D multi-view images and 3D point clouds are powerful, these embeddings are inherently localized to individual objects. Specifically, 3D embeddings capture geometric and spatial attributes of each object in isolation, while 2D embeddings, though derived from full images, are ultimately aggregated per object and offer only limited relational context, making their relational awareness implicit and limited.
As a result, the visual representation lacks explicit modeling of inter-object interactions, limiting its capacity for comprehensive scene understanding and high-level reasoning. To overcome this limitation, we introduce a novel textual modality that encodes relational information in natural language. These descriptions explicitly capture how each object is situated within the scene and how it interacts with others, injecting rich contextual knowledge that is often inaccessible to vision-based embeddings alone.
Our approach introduces this modality through a dual-level integration: (1) by encoding and fusing the text descriptions into the object representations and (2) by injecting them into the prompt space of the language model for enhanced reasoning. This design allows our model to take advantage of both low-level visual embeddings and high-level relational cues, ultimately improving performance on tasks that require contextual understanding.
@@
To encode the generated relational descriptions into the model, we use a Sentence-Transformer model [38] to convert each object’s description into a fixed-dimensional embedding vector $\mathbf{Z}^{v_{t}}_{i}$. This representation is subsequently fused into the object representation, complementing the 2D visual embeddings $\mathbf{Z}^{v_{f}}_{i}$ and the 3D geometric embeddings $\mathbf{Z}^{v_{p}}_{i}$.
@@
In addition to token-level fusion, we leverage a prompt-level strategy that delivers relational priors directly to the language model. When a query mentions a specific object by name or identifier, we prepend the description of that object to the input prompt. To avoid semantic ambiguity and ensure consistent grounding, we replace all object names within the descriptions with their corresponding object identifiers. These object identifiers are learned alongside the language model during training, allowing the LLM to associate them with specific objects in the scene. This strategy eliminates confusion caused by duplicate object names (for example, multiple ’chairs’) and ensures that relational signals remain precisely anchored to the intended objects. The prompt-level text injection allows the LLM to access relevant contextual cues before reading the user’s question, improving reasoning performance.
@@ 3.5 Token Construction and Training Strategy
where $k$ is the length of the response, $s_{i}^{\text{res}}$ is the generated target response, and $s_{[1,\dots,i-1]}^{\text{res}}$ is the previous $i-1$ tokens in the response. $\theta$ denotes the trainable parameters.
@@ 4 Experiments
4.1 Datasets and Metrics
Datasets.
To evaluate the generalizability and task effectiveness of our approach, we conduct experiments on five benchmark datasets spanning four major 3D vision-language tasks. These include: ScanRefer [4] for single-object localization via visual grounding, Multi3DRefer [58] for compositional grounding of multiple targets, Scan2Cap [11] for 3D-aware caption generation, and both ScanQA [2] and SQA3D [34] for contextual question answering within 3D scenes. All five datasets are derived from the ScanNet dataset [13], a richly annotated collection of 1,513 indoor scenes featuring 3D point clouds, RGB images, and camera poses. This shared backbone ensures consistent scene semantics across benchmarks, facilitating multi-task learning. Each benchmark has been aligned to a unified format suitable for instruction-following scenarios. These datasets collectively evaluate a variety of 3D reasoning abilities, ranging from precise localization to compositional understanding.
Metrics.
We use standard evaluation protocols from prior work [22]. For ScanRefer [4], we report thresholded accuracies Acc@0.25 and Acc@0.5, which assess whether the predicted object bounding box has an IoU with the ground truth exceeding 0.25 or 0.5. For Multi3DRefer [58], which involves grounding multiple targets, we use the F1 score at IoU thresholds of 0.25 and 0.5. For the captioning task Scan2Cap [11], we adopt CIDEr@0.5 and BLEU-4@0.5, integrating captioning quality with spatial alignment via IoU. For visual question answering, ScanQA [2] is evaluated using CIDEr and BLEU-4, while SQA3D [34] is evaluated using Exact Match (EM) and its refined variant EM-R as proposed in LEO [23].
4.2 Implementation Details
We extract 100 object proposals per scene using the Mask3D [39] segmentation model. For 3D geometry, we use Uni3D [61] to extract point embeddings $\mathbf{Z}^{v_{p}}_{i}$. For 2D appearance, we adopt DINOv2 [36] to extract 1024-dimensional embeddings per image. We project each object’s 3D mask onto multi-view images and average the cropped DINOv2 embeddings per view. These view-level embeddings are then aggregated using a size-weighted average to obtain the final object-level embeddings $\mathbf{Z}^{v_{f}}_{i}$. For text, we utilize the LLaVA-v1.5-7B [31] model to generate descriptions for each object. These object descriptions are then encoded using the all-mpnet-base-v2 model, a SentenceTransformer [38] based on MPNet [40] that captures rich contextual semantics for downstream fusion, resulting in the embeddings denoted as $\mathbf{Z}^{v_{t}}_{i}$. Each modality is projected to the token space using a three-layer MLP.
We use Vicuna-7B-v1.5 [12] as the LLM and fine-tune it with LoRA [21] with a rank of 16. We train for 3 epochs using a batch size of 32, a base learning rate of $5\times 10^{-6}$, and a cosine annealing schedule. Training is conducted on two 80G NVIDIA A100 GPUs and completes in about 24 hours. We observe that training for 2 epochs yields improved performance on the ScanQA dataset, and we adopt this setting in relevant evaluations.
@@ 4.2 Implementation Details
ScanRefer [4] - - - - - - - 37.3 24.3
ScanQA [2] 30.2 10.1 13.1 33.3 64.9 - - - -
SQA3D [34] - - - - - 46.6 - - -
3DJCG [3] - - - - - - - 49.6 37.3
3D-VLP [27] 30.5 11.1 13.5 34.5 67.0 - - 51.4 39.5
M3DRef-CLIP [58] - - - - - - - 51.9 44.7
3D-VisTA [62] - 13.1 13.9 35.7 72.9 48.5 - 50.6 45.5
ConcreteNet [41] - - - - - - - 50.6 46.5
PQ3D [63] 43.0 - 17.8 - 87.8 47.1 - - 51.2
@@
LAMM [52] - 5.8 - - 42.4 - - - 3.4
Chat-3D [44] 29.1 6.4 11.9 28.5 53.2 - - - -
ZSVG3D [54] - - - - - - - 36.4 32.7
3D-LLM [20] 39.3 12.0 14.5 35.7 69.4 - - 30.3 -
LL3DA [8] - 13.5 15.9 37.3 76.8 - - - -
Grounded 3D-LLM [10] - 13.2 - - 75.9 - - 48.6 44.0
LEO [23] - 11.5 16.2 39.3 80.0 50.0 52.4 - -
Scene-LLM [14] 43.6 12.0 16.6 40.0 80.0 54.2 - - -
3DGraphLLM [56] - 12.1 - - 87.6 53.1 - 57.0 51.3
Chat-Scene [22] 43.2 14.3 18.0 41.6 87.7 54.6 57.5 55.5 50.2
Descrip3D (Ours) 44.0 14.5 18.6 43.1 93.7 55.7 58.4 57.2 51.8

### v2

@@ Abstract
Understanding 3D scenes goes beyond simply recognizing objects; it requires reasoning about the spatial and semantic relationships between them. Current 3D scene-language models often struggle with this relational understanding, particularly when visual embeddings alone do not adequately convey the roles and interactions of objects. In this paper, we introduce Descrip3D, a novel and powerful framework that explicitly encodes the relationships between objects using natural language. Unlike previous methods that rely only on 2D and 3D embeddings, Descrip3D enhances each object with a textual description that captures both its intrinsic attributes and contextual relationships. These relational cues are incorporated into the model through a dual-level integration: embedding fusion and prompt-level injection. This allows for unified reasoning across various tasks such as grounding, captioning, and question answering, all without the need for task-specific heads or additional supervision. When evaluated on five benchmark datasets, including ScanRefer, Multi3DRefer, ScanQA, SQA3D, and Scan2Cap, Descrip3D consistently outperforms strong baseline models, demonstrating the effectiveness of language-guided relational representation for understanding complex indoor scenes. Our code and data are publicly available at https://github.com/jintangxue/Descrip3D.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly transformed human-computer interaction by equipping machines with unprecedented capabilities in language understanding, reasoning, and open-ended dialogue. LLMs have achieved strong results in domains such as retrieval, code, and multi-modal learning [hadi2023survey, minaee2024large, zhao2023survey, li2024survey, xue2024bias, wang2023overview]. Building on this progress, researchers have begun to extend LLMs to 3D scene understanding [ma2024llms, zha2025enable, qi2025gpt4scene], aiming to empower models with the ability to interpret and reason about complex visual and spatial contexts within real-world environments. In particular, indoor scenes are dense and ambiguous. While early models [chen2020scanrefer, chen2021scan2cap] targeted single tasks and recent LLMs [hong20233d, huang2024chat] integrate 2D/3D features, none explicitly model inter-object relations.
@@
This dual-level integration, at both the representation and the language interface levels, allows the model to leverage the prior knowledge of the LLM about the spatial and functional relations learned from the language. Our approach requires no architectural changes, task-specific heads, or extra annotations, making it modular and easy to integrate.
Empirically, our method yields consistent gains across five standard 3D scene language benchmarks. ScanRefer [chen2020scanrefer], Multi3DRefer [zhang2023multi3drefer], Scan2Cap [chen2021scan2cap], ScanQA [azuma2022scanqa], and SQA3D [ma2022sqa3d]. Improvements are especially pronounced in tasks involving complex grounding or multi-object reasoning, highlighting the benefit of explicit text-based relational cues in enhancing 3D scene understanding. Here, multi-object reasoning does not refer only to Multi3DRefer’s multi-target grounding protocol. We use the term more generally to mean queries whose solution depends on object–object relations (e.g., left of, behind, next to, supporting, inside), regardless of whether the benchmark requires multiple boxes.
@@ 2 Related Work
2.1 LLMs for 3D Scene Understanding
Research in 3D vision-language has progressed from task-specific solutions to unified and LLM-driven frameworks. Earlier work focused on single-task models for visual grounding [liu2024survey], dense captioning [yu2023comprehensive], and question answering [li2025embodied]. For example, ScanRefer [chen2020scanrefer], ReferIt3D [achlioptas2020referit3d], and Multi3DRefer [zhang2023multi3drefer] tackled visual grounding by linking natural language to 3D objects. For dense captioning, 3DVG-Transformer [zhao20213dvg] and Vote2Cap-DETR++ [chen2024vote2cap] localized and described objects, while ScanQA [azuma2022scanqa] and SQA3D [ma2022sqa3d] broadened the scope to open-ended scene-level QA. Building on these, multi-task frameworks such as 3DJCG [cai20223djcg] and D3Net [chen2022d3net] unified grounding and captioning with shared encoders, and large-scale pretraining approaches like 3D-VisTA [zhu20233d] and 3D-VLP [jin2023context] established transferable 3D vision-language representations. Yet most treat language as auxiliary, limiting explicit relational modeling.
The emergence of large language models (LLMs) has introduced stronger reasoning and conversational capabilities, motivating their integration with 3D data. Approaches such as PointLLM [xu2024pointllm], ImageBind-LLM [han2023imagebind], Point-Bind [guo2023point], Ulip [xue2023ulip], and CG3D [hegde2023clip] align point cloud embeddings with text, enabling object-level captioning and QA. However, these mainly operate on isolated objects, lacking mechanisms for modeling interactions within complex scenes. To overcome this, multimodal LLMs (MLLMs) have been extended to scene-level reasoning. 3D-LLM [hong20233d] projects 3D embeddings into language space for basic multimodal reasoning, while Chat-3D [wang2023chat] introduces object-centric prompts and staged training to align 3D objects with dialogue. More recently, Chat-Scene [huang2024chat] incorporates object identifiers and fuses them with 2D and 3D embeddings, achieving strong results in QA and grounding benchmarks. Yet, even these advances fall short of explicitly encoding inter-object relationships, leaving relational and spatial reasoning an open challenge for LLM-based 3D scene understanding.
@@ 2.2 Explicit Object Relationship Representation
Although prior models lack explicit mechanisms for encoding inter-object relationships, recent efforts have begun to address this limitation through graph-based or language-based modeling. ConceptGraph [gu2024conceptgraphs] constructs scene-level concept graphs by aligning 3D object embeddings with textual descriptions using vision-language models, enabling semantic relationship modeling and open-vocabulary generalization. HOV-SG [werby2024hierarchical] advances this direction by generating holistic object-scene graphs from 3D scans, capturing spatial and semantic relations across the scene. However, graph-based representations require long sequences and incur high inference costs. To address this, 3DGraphLLM [zemskova20243dgraphllm] incorporates knowledge graph embeddings into object representations, encoding pairwise relationships to support relational reasoning. Yet, this approach introduces significant combinatorial overhead by enumerating object-object triplets and suffers from error propagation in graph construction. Notably, the addition of explicit graph structures in 3DGraphLLM leads to a performance drop on downstream question answering tasks, suggesting that current graph-based formulations may hinder rather than help relational reasoning in practice. Scene-LLM [fu2024scene] instead leverages LLMs with scene-level prompts to generate global summaries and captions, but lacks fine-grained object-level relational modeling. Thus, explicitly and efficiently representing object relationships in 3D scenes remains an open and challenging problem.
2.3 3D Dataset Construction via 2D VLMs
Recent works leverage 2D vision-language models to expand 3D–text resources. Cap3D [luo2023scalable] generates captions for millions of Objaverse objects by aggregating multi-view 2D renderings with pretrained captioning and alignment models. LiDAR-LLM [yang2025lidar] reformulates outdoor LiDAR understanding as a language modeling task, producing hundreds of thousands of LiDAR–text pairs through a staged pipeline. These efforts primarily focus on large-scale data creation to support 3D–language research. In contrast, Descrip3D does not rely on new annotations; instead, it injects object-level descriptions directly into the model pipeline to strengthen relational reasoning across existing benchmarks.
@@ 3.1 Overview
Our method empowers LLMs to perform precise and context-aware reasoning over complex 3D scenes by augmenting object representations with rich, relational textual descriptions. We introduce a powerful text-based relational modality that explicitly encodes both spatial and semantic relationships between objects. Our framework integrates pre-trained 3D and 2D visual encoders, a text encoder, and an LLM, and represents each object as a fused token embedding that combines geometric structure, visual appearance, and relational context. These relational cues are injected into both the object-level embeddings and directly into the LLM prompt, enabling fine-grained multimodal understanding and interaction. This dual-level integration significantly improves the ability of LLM to reason about object relationships, supporting downstream tasks such as grounding, captioning, and question answering with high accuracy and interpretability. An overview of our pipeline is shown in fig.˜2.
@@ 3.2 Object-Centric Scene Representation
We decompose each scene into discrete object proposals using Mask3D [schult2022mask3d] first, producing a set of segmented object point clouds $\{\mathbf{P}_{1},...,\mathbf{P}_{n}\}$. Each object $\mathbf{P}_{i}\in\mathbb{R}^{m_{i}\times 6}$ includes XYZ coordinates and RGB values. Here, $m_{i}$ refers to the number of points in the proposal of the $i$th object. Each resulting object is treated as a fundamental unit for understanding the scene. To support unambiguous reference and interaction, we adopt the design introduced in Chat-Scene [huang2024chat] to associate each object with a unique learnable identifier token $OBJ_{i}\!>$. Concretely, each identifier token is inserted into the tokenizer vocabulary as a learnable lexical token and optimized end-to-end.
@@ 3.3 Visual Embedding Extraction
For each detected object, we extract geometric and visual embeddings from 3D point clouds and multi-view images. For 3D geometry, we use the pre-trained Uni3D [zhou2023uni3d] encoder, which produces an embedding $\mathbf{Z}^{v_{p}}_{i}\in\mathbb{R}^{1\times d}$ from each segmented point cloud $\mathbf{P}_{i}$, and is known for capturing fine-grained structure across diverse scenes. For 2D appearance, we adopt the pre-trained DINOv2 [oquab2023dinov2] vision transformer to extract embeddings from multi-view images. We project each object’s 3D mask onto multi-view images and average the cropped DINOv2 embeddings per view. The features from different views are then aggregated using a size-weighted average into a 2D embedding $\mathbf{Z}^{v_{f}}_{i}\in\mathbb{R}^{1\times d}$. DINOv2 is chosen for its ability to extract semantically rich embeddings from high-resolution input.
@@ 3.4 Text-Based Relational Modality
While the embeddings extracted from both 2D multi-view images and 3D point clouds are powerful, they are inherently localized to individual objects. 3D embeddings capture geometric and spatial attributes in isolation, while 2D embeddings, though derived from full images, are aggregated per object and provide only limited relational context, making their relational awareness implicit and limited.
As a result, the visual representation lacks explicit modeling of inter-object interactions, limiting its capacity for comprehensive scene understanding and high-level reasoning. To overcome this limitation, we introduce a novel textual modality that encodes relational information in natural language. These descriptions capture how each object is situated within the scene and how it interacts with others, injecting contextual knowledge often inaccessible to vision-based embeddings alone.
Our approach integrates this modality through a dual-level integration: (1) by encoding and fusing text descriptions into object representations and (2) by injecting them into the prompt space of the language model for enhanced reasoning. This design allows our model to exploit both low-level visual embeddings and high-level relational cues, ultimately improving performance on tasks requiring contextual understanding.
@@
To encode the generated relational descriptions into the model, we use a Sentence-Transformer model [reimers-2019-sentence-bert] to convert each object’s description into a fixed-dimensional embedding vector $\mathbf{Z}^{v_{t}}_{i}$. This representation is subsequently fused into the object representation, complementing the 2D visual embeddings $\mathbf{Z}^{v_{f}}_{i}$ and the 3D geometric embeddings $\mathbf{Z}^{v_{p}}_{i}$.
@@
In addition to token-level fusion, we leverage a prompt-level strategy that delivers relational priors directly to the language model. When a query mentions an object by name and contains a relational cue (e.g., “next to”), we append a brief description of the relevant object to the end of the input prompt. To avoid semantic ambiguity and ensure consistent grounding, we replace all object names within the descriptions with their corresponding object identifiers. These object identifiers are learned alongside the language model during training, allowing the LLM to associate them with specific objects in the scene. This strategy eliminates confusion caused by duplicate object names (for example, multiple ’chairs’) and ensures that relational signals remain precisely anchored to the intended objects. The prompt-level text injection allows the LLM to access relevant contextual cues before reading the user’s question, improving reasoning performance.
@@ 3.5 Token Construction and Training Strategy
where $k$ is the length of the response, $s_{i}^{\text{res}}$ is the generated target response, and $s_{[1,\dots,i-1]}^{\text{res}}$ is the previous i1i-1 tokens in the response. $\theta$ denotes the trainable parameters.
@@ 4 Experiments
ScanRefer [chen2020scanrefer] - - - - - - - 37.3 24.3
ScanQA [azuma2022scanqa] 30.2 10.1 13.1 33.3 64.9 - - - -
SQA3D [ma2022sqa3d] - - - - - 46.6 - - -
3DJCG [cai20223djcg] - - - - - - - 49.6 37.3
3D-VLP [jin2023context] 30.5 11.1 13.5 34.5 67.0 - - 51.4 39.5
M3DRef-CLIP [zhang2023multi3drefer] - - - - - - - 51.9 44.7
3D-VisTA [zhu20233d] - 13.1 13.9 35.7 72.9 48.5 - 50.6 45.5
ConcreteNet [unal2023three] - - - - - - - 50.6 46.5
PQ3D [zhu2024unifying] 43.0 - 17.8 - 87.8 47.1 - - 51.2
@@
LAMM (Vicuna-13B) [yin2023lamm] - 5.8 - - 42.4 - - - 3.4
Chat-3D (Vicuna-7B) [wang2023chat] 29.1 6.4 11.9 28.5 53.2 - - - -
ZSVG3D (GPT-4) [yuan2024visual] - - - - - - - 36.4 32.7
3D-LLM (BLIP-2-FlanT5) [hong20233d] 39.3 12.0 14.5 35.7 69.4 - - 30.3 -
LL3DA (OPT-1.3B) [chen2024ll3da] - 13.5 15.9 37.3 76.8 - - - -
Grounded 3D-LLM (Tiny-Vicuna-1B) [chen2024grounded] - 13.2 - - 75.9 - - 48.6 44.0
LEO (Vicuna-7B) [huang2023embodied] - 11.5 16.2 39.3 80.0 50.0 52.4 - -
Scene-LLM (LLaMA-2-7B) [fu2024scene] 43.6 12.0 16.6 40.0 80.0 54.2 - - -
3DGraphLLM (Vicuna-7B) [zemskova20243dgraphllm] - 12.1 - - 87.6 53.1 - 57.0 51.3
Chat-Scene (Vicuna-7B) [huang2024chat] 43.2 14.3 18.0 41.6 87.7 54.6 57.5 55.5 50.2
Descrip3D (Ours, Vicuna-7B) 44.0 14.5 18.6 43.1 93.7 55.7 58.4 57.2 51.8